1. The water/cement ratio of a concrete mix is the ratio of the mass of water in the mix to the mass of cement in the mix. If the ratio of a mix is specified as 0.40, how much water should be added to a mixture with 25 kg of cement. 2. The profit y of selling a product is partly constant and party varies directly as the number of products sold x. When x=1, the profit y is $15. When x=100, y=$213. (a) Find the relation between the number of items sold x and the profit y. (b) Find the profit when the number of items sold is 5.
1
Expert's answer
2019-10-23T15:08:44-0400

1) Let "m_w" ​ the mass of water in the mix, "m_c" the mass of cement in the mix.

The ration of water in the mix is "\cfrac{m_w}{m_c}=0.4=\cfrac{m_w}{25 kg}" ​ , "m_w=25*0.4=10kg" .

Answer: 10 kg water should be added to a mixture with 25 kg of cement and the ratio of a mix will be 0.4. 


2) If profit directly depends on sales, then profit can be written through the equation "y=kx+b" , where "x" is number of products sold, and "y" is profit, b is a constant.

If "x=1" then "y=15" ,

If "x=100" then "y=213" .

These values "x,y" must satisfy the equation "y=kx+b" :

"\begin{alignedat}{2} &15 = k+b \\ &213 = 100k+b \end{alignedat}" .

"\begin{alignedat}{2} &15 = k+b \\ &213 = 99k +15 \end{alignedat}" ,

"\begin{alignedat}{2} &15 = k+b \\ &198=99k \end{alignedat}" ,

"\begin{alignedat}{2} &b=13 \\ &k=2 \end{alignedat}".

​So, the relation between "x" and "y" is "y=2x+13" .

If the number of items sold is 5, the profit will be "y = 2*5+13 = 23" .

Answer: a) the ralation is "y=2x+13" ; b) 23.